# Parallel Corpora for Machine Translation in Low-Resource Indic Languages: A Comprehensive Review

### v1

@@ 1.1 Importance of parallel corpora
Parallel corpora are collections of texts that contain sentence-aligned translations across two or more languages  [1]. These resources play a fundamental role in machine translation (MT), cross-lingual natural language processing (NLP), and linguistic research. Unlike monolingual corpora, parallel corpora enable direct learning of translation mappings, making them essential for training statistical and neural MT models  [2].
Parallel corpora have been crucial in the development of MT models, starting from phrase-based statistical MT (SMT)  [3] to modern neural MT (NMT) approaches  [4]. In SMT systems, they provided the necessary data for learning phrase alignments and translation probabilities  [5]. With the rise of transformer-based NMT models  [6], large-scale parallel corpora have become even more critical, as these models rely on extensive aligned data to learn high-quality translation representations.
Beyond MT, parallel corpora are used in cross-lingual NLP tasks such as multilingual word embeddings  [7], zero-shot learning  [8], and multilingual question-answering systems  [9]. These resources allow models to generalize across languages by leveraging shared semantic representations learned from translation pairs.
For low-resource languages, parallel corpora are not just tools for MT but also serve a crucial role in language preservation and revitalization  [10]. Many Indic languages lack digitized linguistic resources, making them vulnerable to digital extinction. Creating high-quality parallel datasets ensures that these languages remain computationally accessible, enabling future educational tools, digital assistants, and automated translations  [11].
Fig. 1: Overview of Machine Translation Model.
@@ 1.2 Machine Translation for Indic languages
MT for Indic languages faces numerous challenges due to their linguistic diversity, script variations, and resource constraints  [12]. Unlike high-resource languages, many Indic languages suffer from limited parallel corpora, making it difficult to train robust translation models. The diversity in syntax and phonology across language families further complicates alignment and translation tasks. The presence of multiple scripts and a lack of standardized transliteration mechanisms hinder effective corpus development.  Figure 3 shows the overview of challenges in Indic MT
Fig. 2: Challenges in Indic Machine Translation: Key issues include morphosyntactic complexity, script variations, low-resource languages, and translation errors in Hinglish and Benglish.
@@ 1.2.1 Morphosyntactic Complexity and Linguistic Variability
Indic languages exhibit significant linguistic diversity, primarily categorized into Indo-Aryan and Dravidian language families  [13]. Indo-Aryan languages, such as Hindi, Bengali, and Marathi, are characterized by inflectional morphology and a relatively flexible subject-object-verb (SOV)  [14] word order, whereas Dravidian languages, including Tamil, Telugu, and Kannada, employ agglutinative morphology, where words are formed by adding multiple affixes to a root word. These structural differences pose challenges in MT systems, as segmentation strategies that work for one language family may not be effective for another  [15]. Additionally, phonological distinctions, such as retroflex consonants in Dravidian languages that are absent in many Indo-Aryan languages, complicate speech-to-text and transliteration tasks  [16].
@@ 1.2.2 Multiscript Representation and Orthographic Challenges
Indic languages are written in multiple scripts, which significantly impact corpus creation and text normalization  [17],  [18]. For instance, Hindi and Marathi share the Devanagari script, but differences in spelling conventions and phonetic representations require preprocessing before effective alignment  [18]. Bengali and Assamese use the Bengali-Assamese script, while Tamil, Telugu, and Kannada have distinct scripts with unique grapheme-to-phoneme mappings  [19].We have also created a indic language categorization illustrated in  Table 2. The lack of script standardization introduces inconsistencies in parallel corpora, making text alignment a challenging task. Moreover, the development of optical character recognition (OCR) tools for Indic scripts remains an ongoing challenge, as many scripts have complex ligatures and diacritic variations that reduce OCR accuracy, further limiting the availability of digitized resources for MT  [20].These script-specific challenges underscore the need for robust preprocessing pipelines and script-aware normalization techniques to improve the quality and usability of Indic language corpora.
@@ 1.2.3 Data Scarcity and Low-Resource Limitations
The scarcity of high-quality parallel corpora remains a significant obstacle in developing robust MT models for Indic languages  [21]. While languages like Hindi and Bengali have relatively larger corpora, low-resource languages such as Santali, Maithili, and Konkani lack sufficient parallel data, restricting the effectiveness of data-driven MT approaches. The limited availability of bilingual datasets hampers the training of neural MT models, which require vast amounts of parallel text for effective generalization. To mitigate this issue, researchers have explored synthetic data generation techniques such as back-translation and cross-lingual transfer learning. However, these approaches often introduce artifacts that can degrade translation quality, highlighting the need for well-annotated, human-verified corpora to support low-resource Indic language MT  [20].
@@ 1.2.4 Register Variability and Linguistic Formality
Most existing parallel corpora for Indic languages are derived from formal sources such as news articles, religious texts, and government documents, which do not capture the informal and conversational aspects of language used in everyday communication  [22]. This imbalance affects the performance of MT systems in real-world applications, as they struggle to translate colloquial expressions, dialectal variations, and code-switched text commonly found in social media and user-generated content  [23]. Code-mixing  [24], particularly in Hindi-English and Bengali-English, presents additional challenges, as standard MT models are not optimized for handling intra-sentential language switching  [25]. The need for diverse corpora that encompass both formal and informal registers is essential to improve translation accuracy across different linguistic contexts.
@@ 2.1 Text-to-Text Parallel Corpora
Text-to-text parallel corpora form the backbone of MT systems, consisting of bilingual or multilingual sentence-aligned datasets that provide direct translations between languages  [26]. These corpora are essential for training supervised MT models and are widely used in both statistical and neural machine translation (NMT) frameworks  [27].
@@ 2.1.1 High-Coverage Parallel Corpora
Large-scale parallel corpora play an important role in developing machine translation systems, especially for low-resource languages. We have considered a dataset to be high-coverage or large-scale if it contains more than 10 million sentence pairs, as this volume provides sufficient linguistic diversity and contextual richness for training robust translation models. These corpora serve as the backbone for both statistical and NMT systems, enabling improved generalization, domain adaptation, and cross-lingual transfer learning. BPCC Parallel Corpus  [28] stands out as the largest, containing 230 million sentence pairs across 22 Indic languages. Its extensive coverage makes it an invaluable resource for multilingual translation tasks, particularly for high-quality English-Indic translations. In comparison, the Samanantar Parallel Corpus  [29], introduced in 2021, includes 46 million sentence pairs between English and 11 Indic languages, along with an additional 82 million sentence pairs between Indic languages. While smaller than BPCC, Samanantar is unique in its extensive Indic-Indic translation pairs, making it highly valuable for intra-Indic translation tasks.
@@
Another notable dataset, CCAligned  [30], consists of over 100 million document pairs across 137 languages, making it one of the most extensive cross-lingual resources. However, this dataset requires extensive filtering to improve data quality before being used for training MT models. Despite its noisiness, its unparalleled scale and diversity make it useful for large-scale pretraining and domain adaptation. The OPUS corpus  [31], widely recognized as a comprehensive multilingual dataset, aggregates multiple parallel corpora across various domains and languages. It covers over 100 million sentence pairs across more than 50 languages, including Hindi, Bengali, Marathi, Tamil, Telugu, and Malayalam, and overlaps significantly with the WAT 2018 Parallel Corpus  [32], which contains 10–20 million sentence pairs specifically for Hindi-English and Bengali-English translations. While OPUS provides domain diversity and structured data, it primarily consists of pre-existing datasets, making it less novel compared to BPCC and Samanantar.
@@
Beyond these major corpora, other high-coverage Indic datasets contribute significantly to MT research. The Bhasha Parallel Corpus  [33] includes 44 million sentence pairs across seven Indic languages, supporting cross-lingual and domain adaptation studies. Additionally, the M2M-100 dataset  [34], developed by Meta AI, contains 12 million sentence pairs spanning over 100 languages, including several Indic languages. This dataset played a key role in the development of the M2M-100 translation model, which enables direct translation between non-English languages.
@@
Several other corpora derived from Wikipedia and Common Crawl have also contributed to large-scale MT training. The WikiMatrix Corpus  [35], developed by Meta, consists of parallel sentences extracted from Wikipedia using LASER-based sentence alignment  [8]. It offers a vast number of sentence pairs across numerous languages, it includes around 3.5 million Hindi-English sentence pairs and around 7 millions of other indic language pairs, making it a valuable resource for training MT models  [36]. However, its reliance on Wikipedia content means that the domain of the text is primarily encyclopedic, limiting its applicability for more conversational or domain-specific translations. Similarly, Wikititles  [37], another Wikipedia-based corpus, extracts bilingual and multilingual article titles from Wikipedia, For Hindi-English, the dataset contains approximately 1.3 million parallel titles. Compared to WikiMatrix, Wikititles provides smaller, well-aligned phrase pairs rather than full sentences, making it particularly useful for training models that focus on short-form content, such as entity names, search queries, or phrase-based MT systems. Due to its structure, Wikititles is less prone to misalignment errors than WikiMatrix but lacks sentence-level parallelism. In contrast, CCMatrix  [35], developed by Meta, is a much larger dataset mined from the CommonCrawl web corpus  [38]. It contains billions of parallel sentences across multiple languages, surpassing both WikiMatrix and Wikititles in sheer volume, making it a powerful resource for large-scale NMT training. However, its primary drawback is the noisiness of web-mined content, which often includes misaligned or irrelevant text pairs that require extensive filtering. Among these datasets, WikiMatrix offers a balanced trade-off between scale and accuracy, providing a large yet relatively clean dataset for training MT models, whereas Wikititles ensures precise alignment quality but is limited in scope due to its focus on article titles rather than full sentences. CCMatrix, on the other hand, offers the most extensive collection of parallel sentences but requires aggressive filtering to ensure usability. While BPCC holds the advantage in terms of sheer size and multilingual support, Samanantar’s inclusion of Indic-Indic translation pairs makes it particularly valuable for intra-Indic translation tasks. Researchers must carefully analyze these corpora to determine the most suitable dataset for their translation models, ensuring an optimal balance between high-quality curated translations and large-scale mined data. The choice of corpus ultimately depends on the specific needs of an MT system—whether prioritizing size, quality, or domain coverage.
@@ 2.2 Low-Coverage Parallel Corpora
The IIT Bombay Parallel Corpus  [39] is one of the most widely used low-resource datasets, containing 1.5 million sentence pairs for English-Hindi translation. The dataset is derived from news and government documents, making it well-suited for formal text translation but less effective for conversational and domain-specific tasks. Due to its clean alignment and high-quality translations, it is often used for benchmarking and fine-tuning NMT models for Hindi-English translation. For Bangla-English translation, the BUET English-Bangla Parallel Corpus  [40] offers 2.7 million sentence pairs, primarily sourced from news articles, books, and religious texts. While smaller than large-scale corpora like BPCC or OPUS, BUET  [40] is an important resource for Bangla machine translation, particularly for formal and literary domains. The dataset provides high-quality bilingual sentence alignments, making it a valuable resource for training MT models that need precise and domain-specific translations.
For historical and classical language translation, the Itihasa Parallel Corpus  [41] is one of the few available datasets, offering 93,000 sentence pairs for English-Sanskrit translation. Given Sanskrit’s morphologically rich structure and complex syntax, this dataset provides a rare opportunity for training translation models in ancient and scholarly texts. Due to its small size, MT models trained on Itihasa rely on data augmentation techniques, such as back-translation and transfer learning, to improve performance. A major initiative for low-resource machine translation across multiple Indic languages is the TICO-19 dataset  [11]. TICO-19 provides parallel sentence pairs across multiple underrepresented Indic languages, including Maithili, Manipuri, and Sindhi. Unlike many general-purpose corpora, TICO-19 is specifically designed for medical and technical translations, making it highly valuable for domain-specific machine translation models. Given the importance of healthcare communication in multilingual settings, this dataset plays a critical role in enabling low-resource language translation for public health applications.
Another key dataset is NLLB (No Language Left Behind)  [42], which provides small-scale training data for multiple Indic languages, including Kashmiri, Maithili, and Bhojpuri. The NLLB project is part of Facebook AI’s initiative to support low-resource language translation, aiming to improve direct translation between non-English language pairs. While individual language pairs in NLLB have limited sentence pairs, the dataset’s wide coverage across many underrepresented Indic languages makes it highly useful for zero-shot and few-shot learning applications in MT.

### v2

@@ Abstract
Parallel Corpora for Machine Translation in Low-Resource Indic Languages: A Comprehensive Review
Rahul Raja Carnegie Mellon University Stanford University LinkedInthanks: Work does not relate to position at LinkedIn.                        Arpita Vats Boston University Santa Clara University LinkedIn11footnotemark: 1
@@ 1.1 Importance of parallel corpora
Parallel corpora are collections of texts that contain sentence-aligned translations across two or more languages  Brown et al. (1991). These resources play a fundamental role in machine translation (MT), cross-lingual natural language processing (NLP), and linguistic research. Unlike monolingual corpora, parallel corpora enable direct learning of translation mappings, making them essential for training statistical and neural MT models  Koehn et al. (2020).
Parallel corpora have been crucial in the development of MT models, starting from phrase-based statistical MT (SMT)  Romdhane et al. (2014) to modern neural MT (NMT) approaches  Stahlberg (2020). In SMT systems, they provided the necessary data for learning phrase alignments and translation probabilities  "Voita and Sennrich . With the rise of transformer-based NMT models  Vaswani et al. (2023), large-scale parallel corpora have become even more critical, as these models rely on extensive aligned data to learn high-quality translation representations.
Beyond MT, parallel corpora are used in cross-lingual NLP tasks such as multilingual word embeddings  Conneau et al. (2020), zero-shot learning  Artetxe and Schwenk (2019), and multilingual question-answering systems  Hu et al. (2024b). These resources allow models to generalize across languages by leveraging shared semantic representations learned from translation pairs.
For low-resource languages, parallel corpora are not just tools for MT but also serve a crucial role in language preservation and revitalization  Hu et al. (2024a). Many Indic languages lack digitized linguistic resources, making them vulnerable to digital extinction. Creating high-quality parallel datasets ensures that these languages remain computationally accessible, enabling future educational tools, digital assistants, and automated translations  Anastasopoulos et al. (2020).
Figure 1: Overview of Machine Translation Model.
@@ 1.2 Machine Translation for Indic languages
MT for Indic languages faces numerous challenges due to their linguistic diversity, script variations, and resource constraints  Bala Das et al. (2023a). Unlike high-resource languages, many Indic languages suffer from limited parallel corpora, making it difficult to train robust translation models. The diversity in syntax and phonology across language families further complicates alignment and translation tasks. The presence of multiple scripts and a lack of standardized transliteration mechanisms hinder effective corpus development.  Figure 2 shows the overview of challenges in Indic MT
Figure 2: Challenges in Indic Machine Translation: Key issues include morphosyntactic complexity, script variations, low-resource languages, and translation errors in Hinglish and Benglish.
@@ 1.2.1 Morphosyntactic Complexity and Linguistic Variability
Indic languages exhibit significant linguistic diversity, primarily categorized into Indo-Aryan and Dravidian language families  Masica (1993). Indo-Aryan languages, such as Hindi, Bengali, and Marathi, are characterized by inflectional morphology and a relatively flexible subject-object-verb (SOV)  Schouwstra and de Swart (2014) word order, whereas Dravidian languages, including Tamil, Telugu, and Kannada, employ agglutinative morphology, where words are formed by adding multiple affixes to a root word. These structural differences pose challenges in MT systems, as segmentation strategies that work for one language family may not be effective for another  Rama and Kolachina (2012). Additionally, phonological distinctions, such as retroflex consonants in Dravidian languages that are absent in many Indo-Aryan languages, complicate speech-to-text and transliteration tasks  Annamalai (2006).
@@ 1.2.2 Multiscript Representation and Orthographic Challenges
Indic languages are written in multiple scripts, which significantly impact corpus creation and text normalization  Manohar et al. (2024),  Hellwig (2010). For instance, Hindi and Marathi share the Devanagari script, but differences in spelling conventions and phonetic representations require preprocessing before effective alignment  Hellwig (2010). Bengali and Assamese use the Bengali-Assamese script, while Tamil, Telugu, and Kannada have distinct scripts with unique grapheme-to-phoneme mappings  Gales et al. (2007).We have also created an indic language categorization illustrated in  Table 2. The lack of script standardization introduces inconsistencies in parallel corpora, making text alignment a challenging task. Moreover, the development of optical character recognition (OCR) tools for Indic scripts remains an ongoing challenge, as many scripts have complex ligatures and diacritic variations that reduce OCR accuracy, further limiting the availability of digitized resources for MT  Sengupta et al. (2019).These script-specific challenges underscore the need for robust preprocessing pipelines and script-aware normalization techniques to improve the quality and usability of Indic language corpora.
@@ 1.2.3 Data Scarcity and Low-Resource Limitations
The scarcity of high-quality parallel corpora remains a significant obstacle in developing robust MT models for Indic languages  Bala Das et al. (2023b). While languages like Hindi and Bengali have relatively larger corpora, low-resource languages such as Santali, Maithili, and Konkani lack sufficient parallel data, restricting the effectiveness of data-driven MT approaches. The limited availability of bilingual datasets hampers the training of neural MT models, which require vast amounts of parallel text for effective generalization. To mitigate this issue, researchers have explored synthetic data generation techniques such as back-translation and cross-lingual transfer learning. However, these approaches often introduce artifacts that can degrade translation quality, highlighting the need for well-annotated, human-verified corpora to support low-resource Indic language MT  Sengupta et al. (2019).
@@ 1.2.4 Register Variability and Linguistic Formality
Most existing parallel corpora for Indic languages are derived from formal sources such as news articles, religious texts, and government documents, which do not capture the informal and conversational aspects of language used in everyday communication  Post et al. (2012). This imbalance affects the performance of MT systems in real-world applications, as they struggle to translate colloquial expressions, dialectal variations, and code-switched text commonly found in social media and user-generated content  Rijhwani et al. (2020). Code-mixing  Khanuja et al. (2020), particularly in Hindi-English and Bengali-English, presents additional challenges, as standard MT models are not optimized for handling intra-sentential language switching  Pratapa et al. (2018). The need for diverse corpora that encompass both formal and informal registers is essential to improve translation accuracy across different linguistic contexts.
@@ 2.1 Text-to-Text Parallel Corpora
Text-to-text parallel corpora form the backbone of MT systems, consisting of bilingual or multilingual sentence-aligned datasets that provide direct translations between languages  de Gibert et al. (2024). These corpora are essential for training supervised MT models and are widely used in both statistical and neural machine translation (NMT) frameworks  Raunak et al. (2024).
@@ 2.1.1 High-Coverage Parallel Corpora
Large-scale parallel corpora play an important role in developing machine translation systems, especially for low-resource languages. We have considered a dataset to be high-coverage or large-scale if it contains more than 10 million sentence pairs, as this volume provides sufficient linguistic diversity and contextual richness for training robust translation models. These corpora serve as the backbone for both statistical and NMT systems, enabling improved generalization, domain adaptation, and cross-lingual transfer learning. BPCC Parallel Corpus  Gala et al. (2023) stands out as the largest, containing 230 million sentence pairs across 22 Indic languages. Its extensive coverage makes it an invaluable resource for multilingual translation tasks, particularly for high-quality English-Indic translations. In comparison, the Samanantar Parallel Corpus  Ramesh et al. (2023), introduced in 2021, includes 46 million sentence pairs between English and 11 Indic languages, along with an additional 82 million sentence pairs between Indic languages. While smaller than BPCC, Samanantar is unique in its extensive Indic-Indic translation pairs, making it highly valuable for intra-Indic translation tasks.
@@
Another notable dataset, CCAligned  El-Kishky et al. (2020), consists of over 100 million document pairs across 137 languages, making it one of the most extensive cross-lingual resources. However, this dataset requires extensive filtering to improve data quality before being used for training MT models. Despite its noisiness, its unparalleled scale and diversity make it useful for large-scale pretraining and domain adaptation. The OPUS corpus  Tiedemann (2012), widely recognized as a comprehensive multilingual dataset, aggregates multiple parallel corpora across various domains and languages. It covers over 100 million sentence pairs across more than 50 languages, including Hindi, Bengali, Marathi, Tamil, Telugu, and Malayalam, and overlaps significantly with the WAT 2018 Parallel Corpus  Zhang et al. (2020), which contains 10–20 million sentence pairs specifically for Hindi-English and Bengali-English translations. While OPUS provides domain diversity and structured data, it primarily consists of pre-existing datasets, making it less novel compared to BPCC and Samanantar.
@@
Beyond these major corpora, other high-coverage Indic datasets contribute significantly to MT research. The Bhasha Parallel Corpus  Mujadia and Sharma (2025) includes 44 million sentence pairs across seven Indic languages, supporting cross-lingual and domain adaptation studies. Additionally, the M2M-100 dataset  Fan et al. (2020), developed by Meta AI, contains 12 million sentence pairs spanning over 100 languages, including several Indic languages. This dataset played a key role in the development of the M2M-100 translation model, which enables direct translation between non-English languages.
@@
Several other corpora derived from Wikipedia and Common Crawl have also contributed to large-scale MT training. The WikiMatrix Corpus  Schwenk et al. (2021), developed by Meta, consists of parallel sentences extracted from Wikipedia using LASER-based sentence alignment  Artetxe and Schwenk (2019). It offers a vast number of sentence pairs across numerous languages, it includes around 3.5 million Hindi-English sentence pairs and around 7 millions of other indic language pairs, making it a valuable resource for training MT models  Niehues and Waibel (2011). However, its reliance on Wikipedia content means that the domain of the text is primarily encyclopedic, limiting its applicability for more conversational or domain-specific translations. Similarly, Wikititles  Liu et al. (2017), another Wikipedia-based corpus, extracts bilingual and multilingual article titles from Wikipedia, For Hindi-English, the dataset contains approximately 1.3 million parallel titles. Compared to WikiMatrix, Wikititles provides smaller, well-aligned phrase pairs rather than full sentences, making it particularly useful for training models that focus on short-form content, such as entity names, search queries, or phrase-based MT systems. Due to its structure, Wikititles is less prone to misalignment errors than WikiMatrix but lacks sentence-level parallelism. In contrast, CCMatrix  Schwenk et al. (2021), developed by Meta, is a much larger dataset mined from the CommonCrawl web corpus  Panchenko et al. (2018). It contains billions of parallel sentences across multiple languages, surpassing both WikiMatrix and Wikititles in sheer volume, making it a powerful resource for large-scale NMT training. However, its primary drawback is the noisiness of web-mined content, which often includes misaligned or irrelevant text pairs that require extensive filtering. Among these datasets, WikiMatrix offers a balanced trade-off between scale and accuracy, providing a large yet relatively clean dataset for training MT models, whereas Wikititles ensures precise alignment quality but is limited in scope due to its focus on article titles rather than full sentences. CCMatrix, on the other hand, offers the most extensive collection of parallel sentences but requires aggressive filtering to ensure usability. While BPCC holds the advantage in terms of sheer size and multilingual support, Samanantar’s inclusion of Indic-Indic translation pairs makes it particularly valuable for intra-Indic translation tasks. Researchers must carefully analyze these corpora to determine the most suitable dataset for their translation models, ensuring an optimal balance between high-quality curated translations and large-scale mined data. The choice of corpus ultimately depends on the specific needs of an MT system—whether prioritizing size, quality, or domain coverage.
@@ 2.2 Low-Coverage Parallel Corpora
The IIT Bombay Parallel Corpus  Kunchukuttan et al. (2018) is one of the most widely used low-resource datasets, containing 1.5 million sentence pairs for English-Hindi translation. The dataset is derived from news and government documents, making it well-suited for formal text translation but less effective for conversational and domain-specific tasks. Due to its clean alignment and high-quality translations, it is often used for benchmarking and fine-tuning NMT models for Hindi-English translation. For Bangla-English translation, the BUETEnglishBanglaCorpus  Islam et al. (2021) offers 2.7 million sentence pairs, primarily sourced from news articles, books, and religious texts. While smaller than large-scale corpora like BPCC or OPUS, BUET  Islam et al. (2021) is an important resource for Bangla machine translation, particularly for formal and literary domains. The dataset provides high-quality bilingual sentence alignments, making it a valuable resource for training MT models that need precise and domain-specific translations.
For historical and classical language translation, the Itihasa Parallel Corpus  Krishna et al. (2020) is one of the few available datasets, offering 93,000 sentence pairs for English-Sanskrit translation. Given Sanskrit’s morphologically rich structure and complex syntax, this dataset provides a rare opportunity for training translation models in ancient and scholarly texts. Due to its small size, MT models trained on Itihasa rely on data augmentation techniques, such as back-translation and transfer learning, to improve performance. A major initiative for low-resource machine translation across multiple Indic languages is the TICO-19 dataset  Anastasopoulos et al. (2020). TICO-19 provides parallel sentence pairs across multiple underrepresented Indic languages, including Maithili, Manipuri, and Sindhi. Unlike many general-purpose corpora, TICO-19 is specifically designed for medical and technical translations, making it highly valuable for domain-specific machine translation models. Given the importance of healthcare communication in multilingual settings, this dataset plays a critical role in enabling low-resource language translation for public health applications.
Another key dataset is NLLB (No Language Left Behind)  Team et al. (2022b), which provides small-scale training data for multiple Indic languages, including Kashmiri, Maithili, and Bhojpuri. The NLLB project is part of Meta AI’s initiative to support low-resource language translation, aiming to improve direct translation between non-English language pairs. While individual language pairs in NLLB have limited sentence pairs, the dataset’s wide coverage across many underrepresented Indic languages makes it highly useful for zero-shot and few-shot learning applications in MT.